# Probabilistic and Differentiable Wireless Simulation with Geometric Transformers

### v1

@@ 6 Discussion
Acknowledgements. We thank Suresh Sharma, Pim de Haan, and Jens Petersen for fruitful discussions.
@@ Appendix A Geometric algebra
As representation, Wi-GATr uses the projective geometric algebra $\mathbb{G}_{3,0,1}$. Here we summarize key aspects of this algebra and define the canonical embedding of geometric primitives in it. For a precise definition and pedagogical introduction, we refer the reader to Dorst [20].
@@
The geometric product is fully defined by bilinearity, associativity, and the condition that the geometric product of a vector with itself is equal to its norm. The geometric product generally maps between different grades. For instance, the geometric product of two vectors will consist of a scalar, the inner product between the vectors, and a bivector, which is related to the cross-product of $\mathbb{R}^{3}$. In particular, the conventional basis elements of grade $k>1$ are constructed as the geometric product of the vector basis elements $e_{i}$. For instance, $e_{12}=e_{1}e_{2}$ is a basis bivector. From the defining properties of the geometric products it follows that the geometric product between orthogonal basis elements is antisymmetric, $e_{i}e_{j}=-e_{j}e_{i}$. Thus, for a $d$-dimensional basis space, there are (${d\choose k}$ independent basis elements at grade $k$.
Projective geometric algebra. To represent three-dimensional objects including absolute positions, we use a geometric algebra based on a base space with $d=4$, adding a homogeneous coordinate to the 3D space. [^4] A three-dimensional base space is not sufficient to represent absolute positions and translations acting on them in a convenient form. See Dorst [20], Ruhe et al. [47], Brehmer et al. [9] for an in-depth discussion. We use a basis $(e_{0},e_{1},e_{2},e_{3})$ with a metric such that $e_{0}^{2}=0$ and $e_{i}^{2}=1$ for $i=1,2,3$. The multivector space is thus $162^{4}=$ 16-dimensional. This algebra is known as the projective geometric algebra $\mathbb{G}_{3,0,1}$.
@@
For more details, we refer the reader to Tbl. 1 in Brehmer et al. [9], or to Dorst [20].
Appendix B Probabilistic model
@@
Formally, we employ the standard DDPM framework [50] to train a latent variable model $p_{\theta}(\mathbf{x}_{0})=\int p_{\theta}(\mathbf{x}_{0:T})d_{\mathbf{x}_{1:T}}$, where $\mathbf{x}_{0}=\left[rsrp,\mathbf{tx},\mathbf{rx},\mathbf{mesh}\right]$ denotes the joint vector of variables following the dataset distribution $p_{data}(\mathbf{x}_{0})$. In DDPM, the latent variables $\mathbf{x}_{1:T}$ are noisy versions of the original data, defined by a discrete forward noise process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right)$ and $\beta_{i}>0$. We approximate the reverse distribution $q(\mathbf{x}_{t-1}\mathbf{x}_{t})$ with $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\sum_{\hat{\mathbf{x}}_{0}}q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})p_{\theta}(\hat{\mathbf{x}}_{0}|\mathbf{x}_{t},t)$, where $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ is a normal distribution with closed-form parameters [25]. The forward and backward distributions $q$ and $p$ form a variational auto-encoder [32] which can be trained with a variational lower bound loss. Using the above parametrization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, however, allows for a simple approximation of this lower bound by training on an MSE objective $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0}}\left[||f_{\theta}(\mathbf{x}_{t},t)-\mathbf{x}_{0}||^{2}\right]$ which resembles denoising score matching [55].
@@
To parametrize $p_{\theta}(\hat{\mathbf{x}}_{0}|\mathbf{x}_{t},t)$, we pass the raw representation of $\mathbf{x}_{t}$ through the wireless GA tokenizer of Wi-GATr and, additionally, we embed the scalar $t$ through a learned timestep embedding [43]. The embedded timesteps can then be concatenated along the scalar channels in the GA representation in a straightforward manner. Similar to GATr [9], the neural network outputs a prediction in the GA representation, which is subsequently converted to the original latent space. Note that this possibly simplifies the learning problem, as the GA representation is inherently higher dimensional than our diffusion space with the same dimensionality as $\mathbf{x}_{0}$.
@@
Equivariant generative modelling. A diffusion model with an invariant base density and an equivariant denoising network defines an invariant density, but equivariant generative modelling has some subtleties [33]. Because the group of translations is not compact, we cannot define a translation-invariant base density. Previous works have circumvented this issue by performing diffusion in the zero center of gravity subspace of euclidean space [27]. However, we found that directly providing the origin as an additional input to the denoising network also resulted in good performance, at the cost of full E(3) equivariance. We also choose to generate samples in the convention where the $z$-axis represents the direction of gravity and positive $z$ is “up”; we therefore provide this direction of gravity as an additional input to our network.
@@
Masking strategies. To improve the performance of conditional sampling, we randomly sample conditioning masks during training which act as an input to the model, as well as a mask on the loss terms. Namely, we sample masks from a discrete distribution with probabilities $p=(0.2,0.3,0.2,0.3)$ corresponding to masks for unconditional, signal, receiver and mesh prediction respectively. If we denote this distribution over masks as $p(m)$, the modified loss function then reads as $\mathcal{L}=\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}),\mathbf{x}_{t},\mathbf{x}_{0}}\left[||\mathbf{m}\odot f_{\theta}(\mathbf{x}_{t}^{\mathbf{m}},t,\mathbf{m})-\mathbf{m}\odot\mathbf{x}_{0}||^{2}\right]$, where $\mathbf{x}^{\mathbf{m}}_{t}$ is equal to $\mathbf{x}_{0}$ along the masked tokens according to $\mathbf{m}$.
Appendix C Datasets
@@
Wi3R dataset. Based on the layouts of the Wi3Rooms dataset by Orekondy et al. [41], we run simulations for 5000 floor layouts that are split into training (4500), validation (250), and test (250). These validation and test splits thus represent generalization across unseen layouts, transmitter, and receiver locations. From the training set, we keep 10 Rx locations as additional test set to evaluate generalization only across unseen Rx locations. To evaluate the generalization performance, we also introduce an out-of-distribution (OOD) set that features four rooms in each of the 250 floor layouts. In all layouts, the interior walls are made of brick while exterior walls are made of concrete. The The Tx and Rx locations are sampled uniformly within the bounds of the floor layouts (10m $\times$ 5m $\times$ 3m).
WiPTR dataset. Based on the floor layouts in the ProcTHOR-10k dataset for embodied AI research [18], we extract the 3D mesh information including walls, windows, doors, and door frames. The layouts comprise between 1 to 10 rooms and can cover up to 600 m2. We assign 6 different dielectric materials for different groups of objects (see Tbl. 5). The 3D Tx and Rx locations are randomly sampled within the bounds of the layout. The training data comprises 10k floor layouts, while test and validation sets each contain 1k unseen layouts, Tx, and Rx locations. Again, we introduce an OOD validation set with 5 layouts where we manually remove parts of the walls such that two rooms become connected. While the multi-modality in combination with the ProcTHOR dataset enables further research for joint sensing and communication in wireless, our dataset set is also, to the best of our knowledge, the first large-scale 3D wireless indoor datasets suitable for embodied AI research.
@@
Appendix D Experiments
D.1 Predictive modelling
@@
For SEGNN, we use representations of up to $\ell_{\text{max}}=3$, 8 layers, and 128 hidden features. The model has $2.6\cdot 10^{5}$ parameters. We selected these parameters in a scan over all three parameters, within the ranges used in Brandstetter et al. [8].
The PLViT model is based on the approach introduced by Hehn et al. [24]. We employ the same centering and rotation strategy as in the original approach around the Tx. Further, we extend the original approach to 3 dimensions by providing the difference in $z$-direction concatenated with the 2D $x$-$y$-distance as one token. Since training from scratch resulted in poor performance, we finetuned a ViT-B-16 model pretrained on ImageNet and keeping only the red channel. This resulted in a model with $85.4\cdot 10^{7}$ parameters and also required us to use a fixed image size for each dataset that ensures the entire floor layout is visible in the image data.
@@
Inference speed. To quantify the trade-off between inference speed and accuracy of signal prediction, we compare the ray tracing simulation with our machine learning approaches. For this purpose, we evaluate the methods on a single room of the validation set with 2 different Tx locations and two equidistant grids at $z\in\{2.3,0.3\}$ with each 1637 Rx locations. Figure 7 summarizes the average inference times per link with the corresponding standard deviation. While Wireless InSite ($6/3/1$, i.e., 6 reflections/3 transmissions/1 diffraction) represents our method that was used to generate the ground truth data, it is also by far the slowest approach. Note that we only measure the inference speed of Wireless InSite for each Tx individually without the preprocessing of the geometry. By reducing the complexity, e.g., reducing the number of allowed reflections or transmissions, of the ray tracing simulation the inference time can be reduced significantly. For example, the configuration $3/2/1$ shows a significant increase in inference speed, but at the same time we can already see that the simulation results do not match the ground truth anymore. This effect is even more pronounced for the case of Wireless InSite $3/1/1$. Our machine learning solutions outperform all tested configurations of Wireless InSite in terms of inference speed, while at the same time keeping competitive performance in terms of prediction accuracy (MAE) compared to the data generation simulation itself in a simpler configuration setting.
@@
Figure 7: Inference wall time vs signal prediction error per Tx/Rx prediction on the first room of the WiPTR validation set.
D.2 Probabilistic modelling
Experiment setup. For all conditional samples involving $p(F_{u}|F_{k},t,r,h)$, we always choose to set $F_{k}$ to be the floor and ceiling mesh faces only and $F_{u}$ to be the remaining geometry. This amounts to completely predicting the exterior walls, as well as the separating walls/doors of the three rooms, whereas the conditioning on $F_{k}$ acts only as a mean to break equivariance. Since $F$ is always canonicalized in the non-augmented training dataset, this allows for direct comparison of variational lower bounds in Tbl. 3 with the non-equivariant transformer baseline.
Models. For both Wi-GATr and the transformer baseline, we follow similar architecture choices as for the predictive models, using an equal amount of attention layers. To make the models timestep-dependent, we additionally employ a standard learnable timestep embedding commonly used in diffusion transformers [43] and concatenate it to the scalar channel dimension.

### v2

@@ Acknowledgements.
We thank Suresh Sharma, Pim de Haan, Maximilian Arnold, and Jens Petersen for fruitful discussions.
@@ Appendix A Geometric algebra
As representation, Wi-GATr uses the projective geometric algebra $\mathbb{G}_{3,0,1}$. Here we summarize key aspects of this algebra and define the canonical embedding of geometric primitives in it. For a precise definition and pedagogical introduction, we refer the reader to Dorst (2020).
@@
The geometric product is fully defined by bilinearity, associativity, and the condition that the geometric product of a vector with itself is equal to its norm. The geometric product generally maps between different grades. For instance, the geometric product of two vectors will consist of a scalar, the inner product between the vectors, and a bivector, which is related to the cross-product of $\mathbb{R}^{3}$. In particular, the conventional basis elements of grade $k>1$ are constructed as the geometric product of the vector basis elements $e_{i}$. For instance, $e_{12}=e_{1}e_{2}$ is a basis bivector. From the defining properties of the geometric products it follows that the geometric product between orthogonal basis elements is antisymmetric, $e_{i}e_{j}=-e_{j}e_{i}$. Thus, for a $d$-dimensional basis space, there are ${\binom{d}{k}}$ independent basis elements at grade $k$.
Projective geometric algebra. To represent three-dimensional objects including absolute positions, we use a geometric algebra based on a base space with $d=4$, adding a homogeneous coordinate to the 3D space. [^4] A three-dimensional base space is not sufficient to represent absolute positions and translations acting on them in a convenient form. See Dorst (2020); Ruhe et al. (2023); Brehmer et al. (2023) for an in-depth discussion. We use a basis $(e_{0},e_{1},e_{2},e_{3})$ with a metric such that $e_{0}^{2}=0$ and $e_{i}^{2}=1$ for $i=1,2,3$. The multivector space is thus $162^{4}=$ 16-dimensional. This algebra is known as the projective geometric algebra $\mathbb{G}_{3,0,1}$.
@@
For more details, we refer the reader to Tbl. 1 in Brehmer et al. (2023), or to Dorst (2020).
Appendix B Wireless Geometric Algebra Tokenizer
Table 3 outlines how the individual scene parameters are embedded into Geometric Algebra.
3D environment ${\bm{F}}$  Triangular mesh 1 token per mesh face  Mesh face center (point)
 Vertices (points)
 Mesh face plane (oriented plane)
 Material classes  One-hot material emb. (scalars)
Antenna ${\bm{x}}^{\text{tx}}$ / ${\bm{x}}^{\text{rx}}$  Position 1 token per antenna  Position (point)
 Orientation  Orientation (direction)
 Receiving / transmitting  One-hot type embedding (scalars)
 Additional characteristics  Characteristics (scalars)
Channel $h$  Antennas 1 token per link  Tx position (point)
 Rx position (point)
 Tx-Rx vector (direction)
 Received power  Normalized power (scalar)
 Phase, delay, …  Additional data (scalars)
Appendix C Diffusion model
@@
Formally, we employ the standard DDPM framework Song et al. (2021) to train a latent variable model $p_{\theta}(\mathbf{x}_{0})=\int p_{\theta}(\mathbf{x}_{0:T})d_{\mathbf{x}_{1:T}}$, where $\mathbf{x}_{0}=\left[rsrp,\mathbf{tx},\mathbf{rx},\mathbf{mesh}\right]$ denotes the joint vector of variables following the dataset distribution $p_{data}(\mathbf{x}_{0})$. In DDPM, the latent variables $\mathbf{x}_{1:T}$ are noisy versions of the original data, defined by a discrete forward noise process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right)$ and $\beta_{i}>0$. We approximate the reverse distribution $q(\mathbf{x}_{t-1}\mathbf{x}_{t})$ with $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\sum_{\hat{\mathbf{x}}_{0}}q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{x}}_{0})p_{\theta}(\hat{\mathbf{x}}_{0}|\mathbf{x}_{t},t)$, where $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ is a normal distribution with closed-form parameters Ho et al. (2020). The forward and backward distributions $q$ and $p$ form a variational auto-encoder Kingma & Welling (2014) which can be trained with a variational lower bound loss. Using the above parametrization of $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, however, allows for a simple approximation of this lower bound by training on an MSE objective $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0}}\left[||f_{\theta}(\mathbf{x}_{t},t)-\mathbf{x}_{0}||^{2}\right]$ which resembles denoising score matching Vincent (2011).
@@
To parametrize $p_{\theta}(\hat{\mathbf{x}}_{0}|\mathbf{x}_{t},t)$, we pass the raw representation of $\mathbf{x}_{t}$ through the wireless GA tokenizer of Wi-GATr and, additionally, we embed the scalar $t$ through a learned timestep embedding Peebles & Xie (2022). The embedded timesteps can then be concatenated along the scalar channels in the GA representation in a straightforward manner. Similar to GATr Brehmer et al. (2023), the neural network outputs a prediction in the GA representation, which is subsequently converted to the original latent space. Note that this possibly simplifies the learning problem, as the GA representation is inherently higher dimensional than our diffusion space with the same dimensionality as $\mathbf{x}_{0}$.
Related Work. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) are a class of generative models that iteratively invert a noising process. They have become the de-facto standard in image and video generation (Ramesh et al., 2022; Ho et al., 2022). Recently, they have also shown to yield promising results in the generation of spatial and sequential data, such as in planning (Janner et al., 2022) and puzzle solving (Hossieni et al., 2024). Aside from their generative modelling capabilities, diffusion models provide a flexible way for solving inverse problems (Chung et al., 2022; Lugmayr et al., 2022; Gloeckler et al., 2024) through multiplication with an appropriate likelihood term (Sohl-Dickstein et al., 2015). Furthermore, by combining an invariant prior distribution with an equivariant denoising network, one obtains equivariant diffusion models (Köhler et al., 2020). These yield a sampling distribution that assigns equal probability to all symmetry transformations of an object, which can improve performance and data efficiency in symmetry problems like molecule generation (Hoogeboom et al., 2022) and planning (Brehmer et al., 2024). We demonstrate similar benefits in modelling wireless signal propagation.
@@
Equivariant generative modelling. A diffusion model with an invariant base density and an equivariant denoising network defines an invariant density, but equivariant generative modelling has some subtleties Köhler et al. (2020). Because the group of translations is not compact, we cannot define a translation-invariant base density. Previous works have circumvented this issue by performing diffusion in the zero center of gravity subspace of euclidean space Hoogeboom et al. (2022). However, we found that directly providing the origin as an additional input to the denoising network also resulted in good performance, at the cost of full E(3) equivariance. We also choose to generate samples in the convention where the $z$-axis represents the direction of gravity and positive $z$ is “up”; we therefore provide this direction of gravity as an additional input to our network.
Unifying forward prediction and inverse problems as conditional sampling. A diffusion model trained to learn the joint density $p_{\theta}({\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$ does not only allow us to generate unconditional samples of wireless scenes, but also lets us sample from various conditionals: given a partial wireless scene, we can fill in the remaining details, in analogy to how diffusion models for images allow for inpainting. To achieve this, we use the conditional sampling algorithm proposed by Sohl-Dickstein et al. (2015): at each step of the sampling loop, we fix the conditioning variables to their known values before feeding them into the denoising network. This algorithm lets us solve signal prediction (sampling from $p_{\theta}(h|{\bm{F}},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}})$), receiver localization (from $p_{\theta}({\bm{x}}^{\text{rx}}|{\bm{F}},{\bm{x}}^{\text{tx}},h)$), geometry reconstruction (from $p_{\theta}({\bm{F}}_{u}|{\bm{F}}_{k},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$), or any other inference task in wireless scenes. We thus unify “forward” and “inverse” modelling in a single algorithm. Each approach is probabilistic, enabling us to model uncertainties. This is important for inverse problems, where measurements often underspecify the solutions.
@@
Masking strategies. In principle, the unconditional diffusion objective should suffice to enable test-time conditional sampling. In practice, we find that we can improve the conditional sampling performance with two modifications. First, we combine training on the unconditional diffusion objective with conditional diffusion objectives. For the latter, we randomly select tokens to condition on and evaluate the diffusion loss only on the remaining tokens. Second, we provide the conditioning mask as an additional input to the denoising model. We sample masks from a discrete distribution with probabilities $p=(0.2,0.3,0.2,0.3)$ corresponding to masks for unconditional, signal, receiver and mesh prediction respectively. If we denote this distribution over masks as $p(m)$, the modified loss function then reads as $\mathcal{L}=\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m}),\mathbf{x}_{t},\mathbf{x}_{0}}\left[||\mathbf{m}\odot f_{\theta}(\mathbf{x}_{t}^{\mathbf{m}},t,\mathbf{m})-\mathbf{m}\odot\mathbf{x}_{0}||^{2}\right]$, where $\mathbf{x}^{\mathbf{m}}_{t}$ is equal to $\mathbf{x}_{0}$ along the masked tokens according to $\mathbf{m}$.
Appendix D Datasets
@@
Wi3R dataset. Based on the layouts of the Wi3Rooms dataset by Orekondy et al. (2022b), we run simulations for 5000 floor layouts that are split into training (4500), validation (250), and test (250). These validation and test splits thus represent generalization across unseen layouts, transmitter, and receiver locations. From the training set, we keep 10 Rx locations as additional test set to evaluate generalization only across unseen Rx locations. To evaluate the generalization performance, we also introduce an out-of-distribution (OOD) set that features four rooms in each of the 250 floor layouts. In all layouts, the interior walls are made of brick while exterior walls are made of concrete. The The Tx and Rx locations are sampled uniformly within the bounds of the floor layouts (10m $\times$ 5m $\times$ 3m).
WiPTR dataset. Based on the floor layouts in the ProcTHOR-10k dataset for embodied AI research Deitke et al. (2022), we extract the 3D mesh information including walls, windows, doors, and door frames. The layouts comprise between 1 to 10 rooms and can cover up to 600 m2. We assign 6 different dielectric materials for different groups of objects (see Tbl. 5). The 3D Tx and Rx locations are randomly sampled within the bounds of the layout. The training data comprises 10k floor layouts, while test and validation sets each contain 1k unseen layouts, Tx, and Rx locations. Again, we introduce an OOD validation set with 5 layouts where we manually remove parts of the walls such that two rooms become connected. While the multi-modality in combination with the ProcTHOR dataset enables further research for joint sensing and communication in wireless, our dataset set is also, to the best of our knowledge, the first large-scale 3D wireless indoor datasets suitable for embodied AI research.
@@
Appendix E Experiments
E.1 Predictive modelling
@@
For SEGNN, we use representations of up to $\ell_{\text{max}}=3$, 8 layers, and 128 hidden features. The model has $2.6\cdot 10^{5}$ parameters. We selected these parameters in a scan over all three parameters, within the ranges used in Brandstetter et al. (2022b).
The PLViT model is based on the approach introduced by Hehn et al. (2023). We employ the same centering and rotation strategy as in the original approach around the Tx. Further, we extend the original approach to 3 dimensions by providing the difference in $z$-direction concatenated with the 2D $x$-$y$-distance as one token. Since training from scratch resulted in poor performance, we finetuned a ViT-B-16 model pretrained on ImageNet and keeping only the red channel. This resulted in a model with $85.4\cdot 10^{7}$ parameters and also required us to use a fixed image size for each dataset that ensures the entire floor layout is visible in the image data.
Finally, we attempt to compare Wi-GATr also to WiNeRT (Orekondy et al., 2022b), a neural ray tracer. However, this architecture, which was developed to be trained on several measurements on the same floor plan, was not able to achieve useful predictions on our diverse datasets with their focus on generalization across floor plans.
@@
Inference speed. To quantify the trade-off between inference speed and accuracy of signal prediction, we compare the ray tracing simulation with our machine learning approaches. For this purpose, we evaluate the methods on a single room of the validation set with 2 different Tx locations and two equidistant grids at $z\in\{2.3,0.3\}$ with each 1637 Rx locations.
The left panel of Fig. 8 summarizes the average inference times per link with the corresponding standard deviation. While Wireless InSite ($6/3/1$, i.e., 6 reflections/3 transmissions/1 diffraction) represents our method that was used to generate the ground truth data, it is also by far the slowest approach. Note that we only measure the inference speed of Wireless InSite for each Tx individually without the preprocessing of the geometry. By reducing the complexity, e.g., reducing the number of allowed reflections or transmissions, of the ray tracing simulation the inference time can be reduced significantly. For example, the configuration $3/2/1$ shows a significant increase in inference speed, but at the same time we can already see that the simulation results do not match the ground truth anymore. This effect is even more pronounced for the case of Wireless InSite $3/1/1$. As of version 0.19, Sionna RT does not support any transmission and only supports first order diffraction, i. e., the path cannot include other reflection events. Therefore, in indoor scenarios such as WiPTR, where signals propagate through walls, Sionna is inaccurate but very fast due to the simplified model and optimized implementation. In addition, it can improve its accuracy by learning more suitable material parameters to compensate for the model simplicity, yet a fundamental gap remains. This shows that if the model is incomplete, the ground truth parameters (see Tab. 5) are not necessarily the optimal parameters. Our neural surrogates provide a favorable trade-off in terms of inference speed and prediction accuracy.
For larger scenes, the inference time of Wi-GATr grows quadratically in the number of mesh faces, just like for a standard transformer. We show this behaviour in the right panel of Fig. 8. Recently, Suk et al. (2024) have successfully demonstrated possibilities to scale GATr to large mesh sizes.
@@
Figure 8: Left: Inference wall time vs signal prediction error per Tx/Rx prediction on the first room of the WiPTR validation set. Right: Scaling of the inference wall time as a function of the number of tokens (mesh faces, antennas, and links) in the scene for Wi-GATr.
E.2 Probabilistic modelling
Experiment setup. For all conditional samples involving $p({\bm{F}}_{u}|{\bm{F}}_{k},{\bm{x}}^{\text{tx}},{\bm{x}}^{\text{rx}},h)$, we always choose to set ${\bm{F}}_{k}$ to be the floor and ceiling mesh faces only and ${\bm{F}}_{u}$ to be the remaining geometry. This amounts to completely predicting the exterior walls, as well as the separating walls/doors of the three rooms, whereas the conditioning on ${\bm{F}}_{k}$ acts only as a mean to break equivariance. Since ${\bm{F}}$ is always canonicalized in the non-augmented training dataset, this allows for direct comparison of variational lower bounds in Tbl. 2 with the non-equivariant transformer baseline.
Models. For both Wi-GATr and the transformer baseline, we follow similar architecture choices as for the predictive models, using an equal amount of attention layers. To make the models timestep-dependent, we additionally employ a standard learnable timestep embedding commonly used in diffusion transformers Peebles & Xie (2022) and concatenate it to the scalar channel dimension.